The following unadjusted trial balance is prepared at fiscal year-end for Nelson Company. Nelson company uses a perpetual inventory system. It categorizes the following accounts as selling expenses: Depreciation Expense—Store Equipment, Sales Salaries Expense, Rent Expense—Selling Space, Store Supplies Expense, and Advertising Expense. It categorizes the remaining expenses as general and administrative.

NELSON COMPANY
Unadjusted Trial Balance
January 31
Debit Credit
Cash $ 23,950
Merchandise inventory 12,500
Store supplies 5,500
Prepaid insurance 2,500
Store equipment 42,800
Accumulated depreciation—Store equipment $ 16,100
Accounts payable 16,000
Common stock 3,000
Retained earnings 35,000
Dividends 2,300
Sales 116,600
Sales discounts 1,950
Sales returns and allowances 2,200
Cost of goods sold 38,000
Depreciation expense—Store equipment 0
Sales salaries expense 15,750
Office salaries expense 15,750
Insurance expense 0
Rent expense—Selling space 7,000
Rent expense—Office space 7,000
Store supplies expense 0
Advertising expense 9,500
Totals $ 186,700 $ 186,700

Additional Information:

  1. Store supplies still available at fiscal year-end amount to $2,550.
  2. Expired insurance, an administrative expense, is $1,750 for the fiscal year.
  3. Depreciation expense on store equipment, a selling expense, is $1,500 for the fiscal year.
  4. To estimate shrinkage, a physical count of ending merchandise inventory is taken. It shows $10,100 of inventory is still available at fiscal year-end.

Problem 4-5A Part 4

4. Compute the current ratio, acid-test ratio, and gross margin ratio as of January 31. (Round your answers to 2 decimal places.)
Answer #1

Nelson Company

Sl no Formula Calculation
a Current ratio Current assets / Current liabilities

= {23,950+10,100+2,550+(2,500-1,750)} / 16,000

= 37,350/16,000

= 2.33

b Acid test ratio (Cash & Cash Equivalent + Accounts receivable) / Current liabilities

= 23,950 / 16,000

= 1.50

c Gross profit ratio (Net sales - Cost of goods sold)/ Net Sales × 100

= (116,600-1950-2200-38,000) / (116,600-1950-2200)

= 66.21%